- 161: Radiation, Site stn2, Fri 22-Oct-1999 13:33:22 CDT, Cleaning radiometers
At 11:15 the domes of the radiometers at station 1 were cleaned
At ~13:00 the domes of the radiometers at stations 2, 3, and 5 were cleaned.
There were traces of brownish grey dust, especially on station 1 domes
All net radiometers were cleaned.
All fans were running. The fan at station 3 was noisy.
- 135: Radiation, Site stn2, Mon 18-Oct-1999 11:59:54 CDT, PYG.OUT at station 2 questioned and fixed (?)
The pyg.out pile was bad so I went out to look at it.
According to Tom, this problem has happened before.
The voltage going to the multiplexer board was varying
quite a bit. I played with the connector inside the Campbell
which resulted in a good signal. The cables had been reversed
from a previous test so I switched them back to the correct
configuration. Will need to monitor to see if fix works. If
so, we will need to flag this Campbell box for repairs.
- 88: Radiation, Site stn2, Tue 12-Oct-1999 17:42:32 CDT, Rpile.out at Station 2
The Rpile.out data has been intermittent at station 2. At 3:50 pm,
local time, Oct 12, I disconnected and reconnected the cable at the
sensor and this appeared to solve the problem. At that time, I also
bumped the Schwartz out of level by popping the leveling screws out
of their upper seats. I reseated the screws and found that the level
was still okay. At that time I also adjusted the level of the net
radiometer about its N/S axis.
The Rpile.out data soon failed again and at 5:34 I swapped the pile.in
and pile.out cables at both ends. This fixed the problem again.
- 45: TRH, Site stn2, Mon 04-Oct-1999 14:40:36 CDT, fix of T/RH problem?
I switched cables (5m instead of 10m to the T/RH) and the Y-adapter on this
station and appear to have gotten both the T/RH and prop to work. It appears
that the T/RH cannot drive 20m of cable with lots of connections.
In the process, I swapped T/RHs between stn5 and stn2, so ID=10 is now at
stn5 and ID=12 is at stn2.